2. On January​ 2, 2019, Kornis Corporation acquired equipment for $300,000. The estimated life of the equipment is 5 years or 40,000 hours. The estimated residual value is $50,000. What is the balance in Accumulated Depreciation on December​ 31, 2019, if Kornis Corporation uses the double−declining−balance method of​depreciation?

A.$120,000

B. $60,000

C.$50,000

D.$100,000

3. On January​ 2, 2019, Konrad Corporation acquired equipment for $760,000. The estimated life of the equipment is 5 years or 37,000 hours. The estimated residual value is $20,000. If Konrad Corporation uses the units of production method of​ depreciation, what will be the debit to Depreciation Expense for the year ended December​ 31, 2020, assuming that during this​ period, the asset was used 5,000 hours?

A.$100,000

B.$148,000

C.$102,703

D.$ 105,405
Answer #1

2.

Double declining depreciation rate = 100 / 5 *2 = 40%

Accumulated depreciation on December 2019 = $300,000*40% = $120,000

Hence option A is correct.

3.

Units of production depreciation rate = $760,000 - 20,000 / 37,000 = $20 per hour

Debit to depreciation expense for the year ended December 31, 2020 = 5,000*$20 = $100,000

Hence option A is correct.